OpenStudy (anonymous):

A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

OpenStudy (anonymous):

Can we have a picture of the Mapping? It would help tremendously.

OpenStudy (texaschic101):

okay.....PQ is perpendicular to AB Now we will look at the line given..-7x + 3y = -21.5 we need to find the slope, so we will put the equation in y = mx + b form, where the number in the m position is the slope. -7x + 3y = - 21.5 --- add 7x to both sides 3y = 7x - 21.5 -- now divide both sides by 3 y = 7/3x - 21.5/3 the slope (m) is 7/3 However, we are looking for a perpendicular line, so we need the negative reciprocal slope. All that means is " flip " the slope and change the sign. So the slope we need is -3/7. (see how I flipped the slope and changed the sign. Notice that PQ is going through points (7,6)....we will use this. Now we use y = mx + b slope(m) = -3/7 (7,6)....x = 7 and y = 6 now we sub y = mx + b 6 = -3/7(7) + b 6 = -3 + b 6 + 3 = b 9 = b so our equation of the central street PQ is : y = -3/7x + 9 and if you need it in standard form : Ax + By = C y = -3/7x + 9 --- multiply everything by 7 7y = -3x + 9 --- add 3 to both sides 3x + 7y = 9 <== in standard form